OpenStudy (anonymous):

solve for z -3z +8= -1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is that the full question, or ?

OpenStudy (anonymous):

yes, is something wrong?

OpenStudy (anonymous):

z = 9/2 is the answer http://www.cymath.com/answer.php?q=z%20-%203z%20%2B%208%20%3D%20-1

OpenStudy (anonymous):

no, just wondering. you should combine like terms. z - 3z = ?

OpenStudy (anonymous):

whoops did the first step wrong, i meant -1 - 8 = -9. thanks @skyelaaboss !

OpenStudy (anonymous):

No problem(: @silverr